Show that 111111111111111111111111111111111111111111111....1 (91times)is not prime

Grade:11

1 Answers

Arun
6 years ago
 
11111111.....(91 times) is a composite number.
Because it is divisible by 1111111.
 
Proof-
11111.....(91 times) can be expressed as-
 = 1/9 * (1091 - 1)
 = 1/9 * (107 - 1) * x = 1111111 * x
Where x = (107)12 + (107)11.....+ 1